CAVITY-ENHANCED, FREQUENCY-AGILE RAPID SCANNING (FARS) SPECTROSCOPY: MEASURE-MENT PRINCIPLES

\underline{JOSEPH T. HODGES}, DAVID A. LONG, GAR-WING TRUONG, KEVIN O. DOUGLASS, STEPHEN E. MAXWELL, ROGER VAN ZEE, and DAVID F. PLUSQUELLIC, National Institute of Standards and Technology, 100 Bureau Drive, Gaithersburg, MD 20899, USA.

We present the principles of frequency-agile, rapid scanning (FARS) spectroscopy, a new technique for high-bandwidth, cavity-enhanced, laser absorption measurements. This method enables a visible or near-infrared probe laser beam to be frequency tuned over several tens of GHz using a microwave source, a waveguide phase modulator and a filter cavity. For the types of cavity-enhanced methods discussed here, the optical resonator itself is used to select a single sideband of the modulated laser spectrum, obviating the need for a separate filter cavity. FARS offers several important advantages over conventional cw laser tuning methods based on thermal or mechanical methods. These include, high speed tuning with sub-ms switching times, the ability to select arbitrary frequency steps or chirp rates, and the realization of a spectrum detuning axis with sub-kHz level precision. We discuss how FARS can be applied to cavity ring-down spectroscopy and other cavity-enhanced methods to enable rapid and accurate measurements of line parameters and to give noise-equivalent absorption coefficients at the 10^{−12} cm^{−1} Hz^{−1/2} level.